OpenStudy (anonymous):

If 2x+y=10 and A=xy, find the maximum value of A.

OpenStudy (blockcolder):

From the first equation, we have \(y=10-2x\). Then, \(A(x)=x(10-2x)=10x-2x^2\). Now, I assume you know how to find the maximum of this quadratic. :)

OpenStudy (anonymous):

I did it this way but the answer isn't correct! :(

OpenStudy (blockcolder):

The maximum value of A is attained when x=5/2. There, A=25-(25/2)=25/2. You mean to say this isn't correct?

OpenStudy (anonymous):

Your answer is correct. How did you do it? I didnt get it. I did it this way: dA/dx= 10 - 4x Then a second derivative for the the max value : dA ^2/ dx^2 =-4

OpenStudy (anonymous):

And i got a wrong answer.

OpenStudy (blockcolder):

\(\Large \frac{d^2A}{dx^2}=-4\) just implies that whatever critical point you get will yield a maximum. (Second Derivative Test) Now, for said critical point, you need to set \(\Large \frac{dA}{dx}=0\), then chug whatever x you get here into A(x).

OpenStudy (anonymous):

Oook !! I got it! Thanks alot and sorry for asking again and again! :)

OpenStudy (blockcolder):

It's okay. :)

OpenStudy (anonymous):

:)